We cannot control whether our bets win or lose, but what we can control is that we get on them at the best odds possible. This will ensure that when you do win, you get paid out the maximum you could have been for that specific bet. If you back a bet at 2.0 and at kick-off time the best price available is 1.80, you know that, win or lose, you’re going into kick-off on the bet at better odds than anyone who was looking to place it just before the game started. It is bet365 according to 48.9% of our users.
No, smaller bookies that provide worse user experiences and features can still be best price on the market.
Some may offer boosted prices, and others may offer money back as a free bet. You want to squeeze as much value as you can get out of any bet you place, and ensuring you place that bet with the bookie that has the biggest odds (best price) on that specific bet is the simplest way to do that. On football, bet365 Early Payout can sometimes give them the edge in the market. Their prices in the match odds markets are usually very competitive and sometimes even best price, with Early Payout included, which means that if the team you back to win goes two goals up at any point you’re paid out as a winner - something that not every bookie offers and even when they do it is rarely at as generous odds as on bet365. Bookmaker overround is the way that the bookies factor in profit margin, best shown as a percentage.
The overround percentage differs across different bookies and markets. They are essentially making sure that there is enough money staked on each side of a bet to ensure that they turn a profit. For example, in any football match, there is a winner on the Both Teams to Score market every time, it is either ‘Yes’ or ‘No’. The bookie can adjust the odds, based on how much money is being staked on Yes and No, to ensure that they turn a profit regardless of the outcome. The overround is the amount beyond 100% that the implied probabilities add up to if you add them up on a specific market.
Take the Both Teams to Score market for Flamengo v Chelsea as an example. Yes is at 2.0 - implied probability of 50%. No is at 1.80 - implied probability of 55.56%. Which means there’s an overround on this market of 5.56%. The markets with the largest overrounds will be the ones that are more challenging to turn a profit on in the long term. The odds change based on how much money is placed on them, that will be different amounts for each bookie therefore, they are often not exactly the same.
In contrast, if you are consistently placing bets that are 'drifting' (going off at bigger odds than when you placed the bet) it is going to be very difficult to turn a profit in the long term. When it comes to visualising the payout you may get from a bet, we prefer to use decimal odds when betting. It gives a clearer representation of the potential return. If you're considering placing a £10 bet and the odds say 11/8 - would you know how much you would be set to receive if it wins? However, if they say 2.38 (11/8 in decimal), you just need to multiply that by your stake to know what your potential return will be bet in play betting football - in this case £23.80.
These three types of odds say the same thing just in a different way, we prefer to use decimal, but it is about finding which one best suits you and which one you enjoy using more when betting. Fractional odds are also known as British odds and are the more traditional odds format used in the UK. A winning £10 bet at 7/1 would return £80. 7 x your stake (in this case £10) = £70 + your stake being returned (in this case £10) = £80. Decimal odds can be easier to use for trading, they reflect the reciprocal of the probability of the outcome.
For example, 5.0 is equal to a probability of ⅕ - a 20% implied probability. Decimal odds also show close to the exact potential return of a bet if the stake was £1. One good thing from a bettor’s perspective is that the bookies are always competing with each other for who can offer the most value. With this in mind, the best way to stay on top of the value is to have multiple accounts. To use horse racing as an example, some bookies will offer more places for each-way terms than others on a certain race. We prefer pre-match betting as there is less time to conduct a thorough research process during the game.
You can guarantee the best odds on horse racing betting by using one of the following bookmakers: Coral, Betfred, Unibet, Bet Goodwin, LiveScore Bet, betway, Ladbrokes, 888sport, talkSPORT Bet, Boyle Sports, Bet MGM, William Hill, Sky Bet, BetVictor and bet365.
| Bookmaker | Avg. Football Margin | Avg. Tennis Margin | Welcome Bonus |
|---|---|---|---|
| Bet365 | 4.2% | 5.1% | $1000 |
| William Hill | 4.8% | 5.5% | $500 |
| Unibet | 4.5% | 5.3% | $750 |
| Betfair | 3.9% | 4.8% | $250 |
Use our free bet calculator to work out your winnings for your bet.
Rule 4 is when a horse is withdrawn from the race, a non-runner, and the odds are adjusted accordingly on the remaining horses in the race. This as less horses being in the race elevates the chance of each remaining horse winning. This will often result in a deduction to the potential returns of the bet you made before the non-runner/s were confirmed. Each-way terms are usually quite similar across all bookies but do differ slightly. Some will offer more places on certain races and may sometimes pay out at a larger fraction of the odds than others.
You want to look for the largest fraction of the odds that is offered in the each-way payout and the most places. On bet365 there is good flexibility for choosing your own number of places on a certain race for an each-way bet. Non-runner no bet means you’ll get your stake back (bet is voided) is the horse is a non-runner. A simple way to begin understanding what value the odds reflect is by converting them to probabilities. Every odds value has an implied probability for example, evens (1/1 or 2.0) has an implied probability of 50%.
With that knowledge and the conducting of a smart research process before every bet, you can start finding bets that your research suggests are of a higher probability than the odds suggest - these are value bets. Another good indicator that you're finding value is the closing price. If you place a bet at a bigger price than it is come kick-off, whether it wins or loses you have found a bit of value. For example, on a Thursday you back Crewe Alexandra to beat Gillingham at 1.80, then come Saturday 3pm (kick-off time) you see that the odds have shortened to 1.70 on the Crewe win. Finding these kinds of bets will likely stand you in good stead when it comes to your profit and loss in the long term. Whether it’s a single or a complex multiple, the Racing Post bet calculator will work out how much you stand to win depending on the eventual outcome. Betting odds are calculated by the probability of something happening.
| Benefit | Impact | Example | Estimated Value Gain |
|---|---|---|---|
| Best Price Discovery | Higher potential payout | 2.10 vs 2.00 odds | +5% per bet |
| Arbitrage Opportunity ID | Guaranteed profit | Back all outcomes | 1-3% profit |
| Bonus & Promotion Tracking | Reduced risk, extra value | Matched bet with bonus | +$50 risk-free |
| Time Saving | Efficient research | Check 20 books in 1 min | Hours per week |
For example, if the odds are 4-1 this suggests there is a one in five chance of winning (4+1), or calculated as 1 / (4+1) = 0.2, which means there is a 20 per cent chance of the outcome happening.
| Tool Name | Key Feature | Sports Covered | Mobile App |
|---|---|---|---|
| Oddschecker | Live Odds Updates | 15+ | Yes |
| OddsPortal | Historical Odds Data | 40+ | Yes |
| BetBrain | Arbitrage Finder | 20+ | No |
| BetExplorer | Form & Statistics | 30+ | Yes |
| SoccerVista | Soccer-Specific Focus | 1 (Soccer) | No |
The winnings you would receive from a bet is calculated by multiplying your stake by the odds. So a stake of £1 would pay £4 profit, plus your stake back, which is a total return of £5. Calculating your winnings from a single selection is fairly straightforward, but when placing more complex bets or multiples it’s advised you use an odds calculator.
Using the reliable Racing Post bet balculator will allow you to work out your returns for bets such as Lucky 15s, doubles, Yankees, accumulators and much more.
With a bet calculator you can compare different outcomes of the different kinds of bets and how they may affect your profits before you even place your bet. It’s incredibly simple to use the betting calculator to work out your winnings.
As part of our Value Bets, we tipped Tranmere to bet gambling bonus sign up beat Crewe @ 1.90 (9.10). The closing price of this bet was 1.80, with that knowledge, we went into the game knowing that we had spotted some value, by taking the bet at a price that was better than anything available at kick-off. Tranmere won the match 2-0, but even if it hadn’t worked out in our favour, it would’ve been a bet that represented value given that the odds shortened towards our selection after placing it. The biggest value indicator is the closing price. Having a positive CLV (closing line value) shows that you have placed a smart bet. The best idea is to start at the top of the page and work your way down through each field.
For player prop betting on smaller leagues and the EFL, bet365 have a better range of odds and prices than any other bookie - while bet365 Sub On Play On covers lots of competitions. Paddy Power, Betfair and Sky Bet are also very competitive and offer the player to be fouled market. Paddy Power came out as the best bookie for bet builders in our survey with 57.3% of the vote. Coral, Betfred, Unibet, Bet Goodwin, LiveScore Bet, betway, Ladbrokes, 888sport, talkSPORT Bet, Boyle Sports, Bet MGM, William Hill, Sky Bet, BetVictor and bet365 all offer Best Odds Guaranteed. This means that if the starting price of a horse ends up higher than the odds you take, you’ll get paid out at the bigger price should you win.
A fixed odds bet combined with Best Odds Guaranteed puts you in the strongest position. Betting on starting price on a bookie without Best Odds Guaranteed means that you may not get the best odds possible for your bet. Whereas Fixed Odds with Best Odds Guaranteed ensures that you do. On Paddy Power, it depends how many are involved in the dead heat for how a bettor is paid out. If it’s a three-way tie for the Top Goalscorer for example, a third of your stake will be treated as a winner and two-thirds as a loser.
A £10 bet at evens would return £6.66 in that circumstance. One third as a winner at evens - £3.33 returns £6.66. The other two-thirds as losers with £3.33 returning 0 and the other £3.33 returning 0. If it were a two-way tie, then half the stake would be treated as a winner. A four-way tie would see a quarter of the stake paid out as a winner, and so on. Follow these simple steps to get started: Select which odds format you want to use Toggle the ‘each-way’ and ‘rule 4’ buttons accordingly Add the odds for each of your selections You’ll now be able to see your expected return and profit from your bet When placing a bet, you may have the option to bet each-way, which means you will get paid out as a winner if your selection wins, but will also be paid out if your selection fails to win but does finish in the top three or four places.